# Can Large Language Models Improve Spectral Graph Neural Networks?

### v1

@@ 4.1 Estimating Homophily Using LLM
After obtaining the homophily labels for a subset of edges using reasoning-based prompting, we further apply consistency-based prompting by querying each sample five times. If at least three out of five predictions yield the same answer, we consider it the final prediction for that sample. The process is illustrated as follows:
$y_{e}=\left\{\begin{array}[]{ll}1&\text{ if }r_{e}\geq 3\\ 0&\text{ if }r_{e}<3\end{array},\right.$
Once the predictions for each sampled edge are obtained, we compute the predicted edge-level homophily $\hat{h}$ as follows:
$\hat{h}=\frac{|\{(u,v)\in\mathcal{E}_{s}\mid\text{LLM}(u,v)=\text{True}\}|}{|\mathcal{E}_{s}|},$ (5)
where $\mathcal{E}_{s}$ denotes the set of all sampled edges.
4.2 Combination of Homophily with Existing Polynomial Filters
After predicting the homophily ratio using an LLM, we incorporate the estimated value into an existing polynomial spectral GNN. To enable spectral GNNs to leverage the homophily information predicted by Eq. (5), we construct a set of heterophily-aware basis vectors inspired by UniFilter (Huang et al. 2024). Specifically, the angle between each pair of basis vectors is defined as follows:
$\theta=\frac{\pi}{2}(1-\hat{h}).$
Thus, the angle between different heterophily-aware basis vectors $\mathbf{u}_{i}$ and $\mathbf{u}_{j}$ is given by:
@@ 4.2 Combination of Homophily with Existing Polynomial Filters
After obtaining the heterophily basis vectors set $\left\{\mathbf{u}_{0},\mathbf{u}_{1},\cdots,\mathbf{u}_{K}\right\}$, we incorporate them into existing polynomial spectral GNNs, such as GPRGNN (Chien et al. 2021), BernNet (He et al. 2021), and JacobiConv (Wang and Zhang 2022).
@@
Table 2: Dataset Statistics
Dataset Children History Fitness Cornell Texas Washington Wisconsin Cora Pubmed Citeseer
Nodes 76,875 41,551 173,055 191 187 229 265 2,708 19,717 3,186
Edges 1,554,578 358,574 1,773,500 292 310 394 510 10,556 88,648 8,450
Feat 768 768 768 768 768 768 768 1,433 500 3,703
Classes 24 12 13 5 5 5 5 7 3 6
Homophily 0.42 0.66 0.90 0.12 0.06 0.15 0.17 0.81 0.80 0.78
@@ 4.3 Training Objective
After integrating the LLM-estimated homophily ratio into existing polynomial spectral filters, node classification can be performed using various polynomial-based spectral GNNs. Notably, the proposed method introduces no additional trainable parameters; it simply incorporates heterophily-aware basis vectors guided by the estimated homophily.
@@
where $\mathcal{V}_{\text{train }}$ is the training node set, and $\mathbf{y}^{j}$ is the ground-truth one-hot label vector of node $j$.
@@ 5 Experiment
  • RQ1: Can the performance of polynomial spectral GNNs be improved by incorporating homophily estimated by LLMs?
  • RQ2: How does the homophily estimated by LLMs compare to other estimation methods in enhancing polynomial spectral GNNs?
  • RQ3: Does the LLM-based homophily estimation incur lower inference cost?
  • RQ4: What is the impact of the key hyperparameter $\beta$ on model performance?
@@ 5.1 Experimental Setup
Datasets. We select ten graph datasets with text attributes , including three citation networks (Cora, Citeseer, and Pubmed), four webpage networks (Cornell, Texas, Washington, and Wisconsin), and three Amazon co-purchase networks (Children, History, and Fitness). These datasets cover both homophilic and heterophilic graph structures. The statistics of these datasets are summarized in Table 2.
@@
Table 3: Performance (%) on Various Datasets (mean with standard deviation as subscript)
@@
Method Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
MLP 80.85±0.00 84.91±0.00 68.38±1.48 79.82±1.52 54.90±0.04 83.91±0.14 79.57±0.06 79.52±0.32 71.68±1.68 87.43±0.11
GCN 57.45±4.26 35.85±3.77 37.61±10.36 57.02±1.52 56.47±0.10 84.82±0.05 90.86±0.06 86.22±0.57 78.58±0.81 85.02±0.21
GraphSAGE 78.72±2.13 78.62±1.09 69.23±4.44 76.32±2.64 54.39±1.32 84.51±0.19 88.55±2.07 86.65±0.77 76.96±1.36 88.80±0.31
GAT 44.68±11.85 38.36±4.75 30.77±10.26 51.75±15.20 49.78±4.31 82.79±1.18 90.51±0.68 84.93±0.76 78.16±0.45 85.04±0.38
APPNP 68.79±2.45 54.72±0.00 57.26±1.48 62.28±1.52 45.34±0.27 81.42±0.07 80.33±0.36 88.01±0.64 81.40±0.33 87.58±0.03
[…]
UniFilter 88.61±4.16 84.52±3.01 69.00±4.12 79.38±3.65 47.55±0.16 86.99±0.84 87.81±4.74 87.19±1.73 77.57±2.10 85.55±3.63
GPRGNN 82.98±2.13 79.87±4.75 68.38±2.96 81.58±2.63 60.29±0.46 86.03±0.07 92.38±0.07 88.01±0.19 78.16±0.33 89.11±0.19
GPRGNNPLUS 87.23±0.00 81.76±2.18 70.94±2.96 83.33±1.52 61.00±0.20 86.47±0.13 92.58±0.04 88.07±0.21 79.94±0.41 89.53±0.09
BernNet 79.43±1.23 84.91±1.89 67.52±1.48 80.70±1.52 58.76±0.17 84.93±0.30 91.83±0.16 82.41±0.60 72.83±1.48 88.54±0.36
BernNetPLUS 90.78±1.23 85.53±2.18 71.79±5.13 83.33±1.52 60.80±0.19 85.66±0.09 92.44±0.06 87.58±1.06 78.89±1.28 89.37±0.08
JacobiConv 78.01±2.45 84.28±1.09 64.96±3.92 79.82±3.04 59.63±0.46 85.36±0.13 90.70±0.17 84.99±1.01 72.88±1.13 88.04±0.10
JacobiConvPLUS 80.85±0.00 84.91±1.89 68.38±1.48 80.70±1.52 59.67±0.05 85.79±0.07 91.04±0.26 88.25±0.39 78.63±1.48 89.48±0.13
ChebNetII 82.27±1.00 85.53±0.89 68.38±1.21 84.21±2.15 56.74±0.07 84.76±0.10 83.49±0.02 81.01±0.12 72.37±0.24 88.34±0.13
ChebNetIIPLUS 87.94±1.00 85.53±0.89 69.23±2.09 84.21±2.15 58.76±0.12 85.79±0.05 88.22±0.17 83.53±0.17 74.12±0.36 88.34±0.10
@@
Settings. We adopt the experimental setup used in CS-TAG (Yan et al. 2023) and follow its standard data split to ensure a fair comparison. Specifically, the training/validation/test sets are divided as 60%/20%/20% for all datasets except Fitness, for which the split is 20%/10%/70%. We use the accuracy metric as an evaluation indicator. All experiments are performed three times, and we report the average results and their corresponding standard errors. All experiments are conducted on a machine with 3 NVIDIA A5000 24GB GPUs and Intel(R) Xeon(R) Silver 4310 2.10 GHz CPU.
@@
Baselines. To thoroughly evaluate the effectiveness of the proposed method, in addition to the three polynomial spectral GNN backbones—GPR-GNN (Chien et al. 2021), BernNet (He et al. 2021), JacobiConv (Wang and Zhang 2022), and ChebNetII (He, Wei, and Wen 2022)—we also include four classical GNN models and a Multi-Layer Perceptron (MLP) as baselines: GCN (Kipf and Welling 2017), GAT (Veličković et al. 2019), GraphSAGE (Hamilton, Ying, and Leskovec 2017), APPNP (Gasteiger, Bojchevski, and Günnemann 2019), TFE-GNN (Duan et al. 2024), and UniFilter (Huang et al. 2024).
@@ 5.2 Main Results (RQ1)
Table 3 presents the node classification results across ten datasets. As observed, the proposed method consistently outperforms GPRGNN, BernNet, JacobiConv, and ChebNetII by a significant margin. Specifically, it achieves an average improvement of 1.14% over GPRGNN, 4.51% over BernNet, 2.51% over JacobiConv, and 2.41% over ChebNetII. These results clearly demonstrate the effectiveness of the proposed method. Moreover, although recent methods such as ChebNetII and TFE-GNN have shown competitive performance, our method still surpasses them. This highlights the strong representational power of LLM-enhanced spectral GNNs.
@@ 5.3 Ablation Analysis (RQ2)
This subsection aims to evaluate the advantage of using LLM-predicted homophily in polynomial spectral GNNs. We consider the following two alternative variants for comparison:
  1. 1.
    Training a MLP to predict homophily and enhance the spectral GNN.
  2. 2.
    Directly using the homophily computed from the training set to enhance the spectral GNN.
Figure 2 compares the performance of spectral GNNs enhanced with homophily estimated by different methods on four datasets  [^1] Please refer to Appendix B for more results.. As shown, the homophily predicted by LLMs yields the best performance in nearly all cases, with the only exception being BernNet on the Texas dataset. This demonstrates the superiority of LLM-estimated homophily over that derived from MLPs or directly from the training set, likely due to the stronger reasoning capability of LLMs. In addition, MLP-based homophily estimation generally outperforms the one directly obtained from the training set, indicating that learning homophily via a trainable model is more effective than relying on limited labeled data. However, MLP-based variants still fall short of the LLM-based approach, highlighting that stronger models can generate more informative homophily signals, which in turn lead to better performance when integrated into spectral GNNs.
@@
Type Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
Output 0.12M/0.07\$ 0.13M/0.08\$ 0.14M/0.09\$ 0.13M/0.08\$ 0.17M/0.10\$ 0.16M/0.10\$ 0.15M/0.09\$ 0.17M/0.10\$ 0.16M/0.10\$ 0.18M/0.11\$
Input 0.47M/0.07\$ 0.67M/0.10\$ 0.56M/0.08\$ 0.44M/0.07\$ 0.38M/0.06\$ 0.36M/0.05\$ 0.09M/0.01\$ 0.22M/0.03\$ 0.24M/0.04\$ 0.40M/0.06\$
Total 0.59M/0.14\$ 0.80M/0.18\$ 0.71M/0.17\$ 0.57M/0.14\$ 0.55M/0.16\$ 0.53M/0.15\$ 0.24M/0.10\$ 0.39M/0.13\$ 0.40M/0.13\$ 0.58M/0.17\$
@@
Method Washington Wisconsin Cornell Texas Children History Fitness Cora Citeseer Pubmed
GPRGNN 13.00/2.60 12.30/2.46 11.80/2.36 13.40/2.67 230.80/46.16 94.90/18.97 472.70/94.54 15.00/3.00 31.40/6.29 32.60/6.52
GPRGNNPLUS 13.60/2.72 14.00/2.81 13.90/2.77 13.90/2.78 274.00/54.80 119.10/23.83 575.40/115.08 18.80/3.76 41.40/8.28 40.60/8.12
BernNet 37.50/7.50 36.90/7.38 38.60/7.72 35.90/7.18 1235.30/247.06 445.70/89.15 2449.40/489.88 50.20/10.04 132.10/26.42 128.30/25.67
BernNetPLUS 43.50/8.69 40.90/8.17 43.80/8.76 43.50/8.71 1303.50/260.70 473.80/94.77 2596.20/519.23 53.70/10.73 143.10/28.63 138.20/27.65
JacobiConv 26.80/5.36 26.80/5.36 26.80/5.36 26.80/5.35 56.80/11.37 33.90/6.77 88.00/17.60 28.80/5.76 28.10/5.63 30.10/6.02
JacobiConvPLUS 28.10/5.63 28.60/5.72 27.90/5.58 28.00/5.59 56.50/11.31 34.40/6.88 86.30/17.26 30.50/6.09 29.80/5.96 31.30/6.26
ChebNetII 49.40/9.88 50.20/10.05 49.60/9.93 49.20/9.83 274.70/54.95 137.00/27.41 574.50/114.89 51.80/10.35 70.40/14.08 68.80/13.76
ChebNetIIPLUS 50.00/10.01 50.70/10.13 50.60/10.12 51.20/10.24 324.20/64.84 164.50/32.90 685.70/137.13 54.80/10.95 79.50/15.89 76.90/15.38

### v2

@@ 4.2 Incorporating Homophily into Polynomial Spectral Filters
[…]
@@
After obtaining the heterophily basis vectors set $\left\{\mathbf{u}_{0},\mathbf{u}_{1},\cdots,\mathbf{u}_{K}\right\}$, we incorporate them into existing polynomial spectral GNNs, such as GPRGNN (Chien et al. 2021), BernNet (He et al. 2021), JacobiConv (Wang and Zhang 2022), and ChebNetII (He, Wei, and Wen 2022). The details are as follows:
@@
Table 2: Dataset Statistics
Dataset Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
Nodes 229 191 187 76,875 41,551 173,055 2,708 3,186 19,717
Edges 394 292 310 1,554,578 358,574 1,773,500 5,429 4,277 44,335
Feat 768 768 768 768 768 768 1,433 3,703 500
Classes 5 5 5 24 12 13 7 6 3
Homophily 0.15 0.12 0.06 0.42 0.66 0.90 0.81 0.78 0.80
@@
Insertion Details of ChebNetII. To mitigate overfitting and reduce the Runge phenomenon, ChebNetII (He, Wei, and Wen 2022) enhances the original Chebyshev (Defferrard, Bresson, and Vandergheynst 2016) polynomial approximation by reparameterizing the filter response as a learnable coefficient $\gamma_{j}$ , as shown below:
$\mathbf{z}=\frac{2}{K+1}\sum_{k=0}^{K}\sum_{j=0}^{K}\gamma_{j}T_{k}\left(x_{j}\right)T_{k}(\mathbf{L})\mathbf{x},$ (13)
where $x_{j}=\cos((j+1/2)\pi/(K+1))$ are the Chebyshev nodes of $T_{K+1}$. The recursive definition of the Chebyshev polynomial $T_{k}(x)$ is given as follows:
$T_{0}(x)$ $=1,$ (14)
$T_{1}(x)$ $=x,$
$T_{k}(x)$ $=2xT_{k-1}(x)-T_{k-2}(x).$
We incorporate the heterophily basis vector $\mathbf{u}_{k}$ into the corresponding $k$-th term of the ChebNetII, resulting in:
$\mathbf{z}=\frac{2}{K+1}\sum_{k=0}^{K}\left[\beta\sum_{j=0}^{K}\gamma_{j}T_{k}\left(x_{j}\right)T_{k}(\mathbf{L})\mathbf{x}+(1-\beta)\mathbf{u}_{k}\right],$ (15)
@@ 4.3 Training Objective
After integrating the LLM-estimated homophily ratio into existing polynomial spectral filters, node classification can be performed using various polynomial-based spectral GNNs. Notably, the proposed method introduces no additional trainable parameters.
@@
[…]
@@ 5 Experiment
  • RQ1: How does the proposed method perform compared to its corresponding spectral GNN counterparts?
  • RQ2: How effective is LLM-predicted homophily compared to that estimated by MLPs or directly derived from the training set?
  • RQ3: Does the LLM-based homophily estimation incur negligible time and money cost?
  • RQ4: How does the key hyperparameter $\beta$ influence the model’s performance?
@@ 5.1 Experimental Setup
Datasets. We select nine graph datasets with text attributes (Yan et al. 2023; Guo et al. 2024) , including three citation networks (Cora, Citeseer, and Pubmed), three webpage networks (Cornell, Texas, and Washington), and three Amazon co-purchase networks (Children, History, and Fitness). The statistics of these datasets are summarized in Table 2.
@@
Table 3: Performance (%) on Various Datasets (mean accuracy with standard deviation as subscript)
@@
Method Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
MLP 80.85±0.00 68.38±1.48 79.82±1.52 54.90±0.04 83.91±0.14 79.57±0.06 79.52±0.32 71.68±1.68 87.43±0.11
GCN 57.45±4.26 37.61±10.36 57.02±1.52 56.47±0.10 84.82±0.05 90.86±0.06 86.22±0.57 78.58±0.81 85.02±0.21
GraphSAGE 78.72±2.13 69.23±4.44 76.32±2.64 54.39±1.32 84.51±0.19 88.55±2.07 86.65±0.77 76.96±1.36 88.80±0.31
GAT 44.68±11.85 30.77±10.26 51.75±15.20 49.78±4.31 82.79±1.18 90.51±0.68 84.93±0.76 78.16±0.45 85.04±0.38
APPNP 68.79±2.45 57.26±1.48 62.28±1.52 45.34±0.27 81.42±0.07 80.33±0.36 88.01±0.64 81.40±0.33 87.58±0.03
TFEGNN 68.09±9.19 66.67±5.54 64.04±1.24 45.44±0.83 81.19±0.03 83.31±1.02 85.50±0.27 77.83±0.52 88.25±0.17
UniFilter 88.61±4.16 69.00±4.12 79.38±3.65 47.55±0.16 86.99±0.84 87.81±4.74 87.19±1.73 77.57±2.10 85.55±3.63
GPRGNN 82.98±2.13 68.38±2.96 81.58±2.63 60.29±0.46 86.03±0.07 92.38±0.07 88.01±0.19 78.16±0.33 89.11±0.19
GPRGNNPLUS 87.23±0.00 70.94±2.96 83.33±1.52 61.00±0.20 86.47±0.13 92.58±0.04 88.07±0.21 79.94±0.41 89.53±0.09
BernNet 79.43±1.23 67.52±1.48 80.70±1.52 58.76±0.17 84.93±0.30 91.83±0.16 82.41±0.60 72.83±1.48 88.54±0.36
BernNetPLUS 90.78±1.23 71.79±5.13 83.33±1.52 60.80±0.19 85.66±0.09 92.44±0.06 87.58±1.06 78.89±1.28 89.37±0.08
JacobiConv 78.01±2.45 64.96±3.92 79.82±3.04 59.63±0.46 85.36±0.13 90.70±0.17 84.99±1.01 72.88±1.13 88.04±0.10
JacobiConvPLUS 80.85±0.00 68.38±1.48 80.70±1.52 59.67±0.05 85.79±0.07 91.04±0.26 88.25±0.39 78.63±1.48 89.48±0.13
ChebNetII 82.27±1.00 68.38±1.21 84.21±2.15 56.74±0.07 84.76±0.10 83.49±0.02 81.01±0.12 72.37±0.24 88.34±0.13
ChebNetIIPLUS 87.94±1.00 69.23±2.09 84.21±2.15 58.76±0.12 85.79±0.05 88.22±0.17 83.53±0.17 74.12±0.36 88.34±0.10
@@
Settings. We adopt the experimental setup used in CS-TAG (Yan et al. 2023) and follow its standard data split to ensure a fair comparison. Specifically, the training/validation/test sets are divided as 60%/20%/20% for all datasets except Fitness, for which the split is 20%/10%/70%. We use the accuracy metric as an evaluation indicator. All experiments are performed three times, and we report the average results and their corresponding standard errors. All experiments are conducted on a machine with 3 NVIDIA A5000 24GB GPUs and Intel(R) Xeon(R) Silver 4310 2.10 GHz CPU. The LLM used for generating homophily ratio predictions is GPT-4o-mini (Hurst et al. 2024).
@@
Baselines. To thoroughly evaluate the effectiveness of the proposed method, in addition to the four polynomial spectral GNN backbones—GPR-GNN (Chien et al. 2021), BernNet (He et al. 2021), JacobiConv (Wang and Zhang 2022), and ChebNetII (He, Wei, and Wen 2022)—we also include six classical GNN models and a Multi-Layer Perceptron (MLP) as baselines: GCN (Kipf and Welling 2017), GAT (Veličković et al. 2019), GraphSAGE (Hamilton, Ying, and Leskovec 2017), APPNP (Gasteiger, Bojchevski, and Günnemann 2019), TFE-GNN (Duan et al. 2024), and UniFilter (Huang et al. 2024).
For more experimental details, including comprehensive descriptions of the baselines, datasets, and hyperparameter settings, please refer to Appendix A.
@@ 5.2 Main Results (RQ1)
Table 3 reports the node classification accuracies on nine benchmark datasets. As observed, our LLM-enhanced variants (denoted with the PLUS suffix) consistently outperform their original counterparts, including GPRGNN, BernNet, JacobiConv, and ChebNetII, across all datasets. Concretely, the average performance gains over the original models are substantial: +1.17% for GPRGNN, +4.93% for BernNet, +2.70% for JacobiConv, and +2.68% for ChebNetII. These consistent improvements clearly demonstrate the effectiveness of incorporating LLM-predicted homophily into spectral GNNs.
In addition, despite the competitive performance of recent spectral methods such as UniFilter (Huang et al. 2024) and TFE-GNN (Duan et al. 2024), our approach still achieves superior results across diverse datasets. Importantly, our method introduces no graph topology modifications or task-specific training, yet enhances the adaptability of spectral GNNs to different structural patterns, including both homophilic and heterophilic graphs.
Overall, the results demonstrate that incorporating global structural priors via LLMs can significantly boost the representational power of spectral GNNs, offering a simple yet effective enhancement to existing models.
@@ 5.3 Ablation Analysis (RQ2)
We evaluate the impact of incorporating LLM-predicted homophily into spectral GNNs (LLM-based) by comparing it with two variants: (1) MLP-based, which trains an MLP to classify each edge as homophilic or heterophilic for estimating overall homophily; (2) Train-label-based, which directly computes homophily from the proportion of homophilic edges among labeled node pairs in the training set. Figure 2 presents the performance of these variants on four datasets. As shown, LLM-predicted homophily consistently achieves the best results, with the only exception being BernNet on the Texas dataset. This underscores the advantage of LLMs in generating more reliable structural priors, likely due to their stronger reasoning capabilities. Notably, MLP-based estimation also outperforms the direct use of training-set homophily, indicating that learning homophily signals is more effective than relying on limited labeled data. However, it still lags behind the LLM-based variant, highlighting that the quality of the estimated homophily plays a crucial role in guiding spectral filter construction.
@@
Type Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
Input 0.47/0.07\$ 0.56/0.08\$ 0.44/0.07\$ 0.38/0.06\$ 0.36/0.05\$ 0.09/0.01\$ 0.22/0.03\$ 0.24/0.04\$ 0.40/0.06\$
Output 0.12/0.07\$ 0.14/0.09\$ 0.13/0.08\$ 0.17/0.10\$ 0.16/0.10\$ 0.15/0.09\$ 0.17/0.10\$ 0.16/0.10\$ 0.18/0.11\$
Total 0.59/0.14\$ 0.71/0.17\$ 0.57/0.14\$ 0.55/0.16\$ 0.53/0.15\$ 0.24/0.10\$ 0.39/0.13\$ 0.40/0.13\$ 0.58/0.17\$
@@
Method Washington Cornell Texas Children History Fitness Cora Citeseer Pubmed
GPRGNN 13.0/2.6 11.8/2.4 13.4/2.7 230.8/46.2 94.9/19.0 472.7/94.5 15.0/3.0 31.4/6.3 32.6/6.5
GPRGNNPLUS 13.6/2.7 13.9/2.8 13.9/2.8 274.0/54.8 119.1/23.8 575.4/115.1 18.8/3.8 41.4/8.3 40.6/8.1
BernNet 37.5/7.5 38.6/7.7 35.9/7.2 1235.3/247.1 445.7/89.2 2449.4/489.9 50.2/10.0 132.1/26.4 128.3/25.7
BernNetPLUS 43.5/8.7 43.8/8.8 43.5/8.7 1303.5/260.7 473.8/94.8 2596.2/519.2 53.7/10.7 143.1/28.6 138.2/27.7
JacobiConv 26.8/5.4 26.8/5.4 26.8/5.4 56.8/11.4 33.9/6.8 88.0/17.6 28.8/5.8 28.1/5.6 30.1/6.0
JacobiConvPLUS 28.1/5.6 27.9/5.6 28.0/5.6 56.5/11.3 34.4/6.9 86.3/17.3 30.5/6.1 29.8/6.0 31.3/6.3
ChebNetII 49.4/9.9 49.6/9.9 49.2/9.8 274.7/55.0 137.0/27.4 574.5/114.9 51.8/10.4 70.4/14.1 68.8/13.8
ChebNetIIPLUS 50.0/10.0 50.6/10.1 51.2/10.2 324.2/64.8 164.5/32.9 685.7/137.1 54.8/11.0 79.5/15.9 76.9/15.4